OpenStudy (virnios):

-3x - 8y = 10, Find the value of y when x equals 2

OpenStudy (alexandervonhumboldt2):

substitute 2 for x: -3*2-8y=10 -6-8y=10 -8y=16 -8y/-8=16/-8 y=-2